OpenStudy (anonymous):

(cos xy+x/y)dx + (1+x/y cosxy)dy=0 solve the differentional equation

OpenStudy (experimentx):

\[ \left ( \cos (xy)+ { x \over y} \right )dx + \left (1+{x \over y} \cos(xy) \right )dy \]

OpenStudy (experimentx):

??